OpenStudy (anonymous):

Suppose A is (3,2,-1) and B is (-1,2,4) a. Write down the vector equation of the line through A and B b. Find 2 points on the line (AB) which are 2√11 units from A

OpenStudy (kesumonu):

\[\frac{ x-3 }{ -4 }=\frac{ y-2 }{ 0 }=\frac{ z+1 }{ 5}\] this is the equation of line passing through two given points A and B

OpenStudy (kesumonu):

since y coordinate of the two point are the same they are in same plane